Three identical charges are fixed at the corners of an equilateral triangle  If the force between any two charges is F then the net electric force on any one of them 
1)  root 2F
2)  2F
3)  root 3 F
4)  3F

Grade:12

2 Answers

Piyush Kumar Behera
4 years ago
This is a vector addition question.As each of the particle is exering an repulsive force on each other of magnitude F.
As the triangle given is an equilateral triangle. Hence the angle between the two forces is 60 degrees.
 
So net force = \sqrt{F^{2}+F^{2}+2F^{2}cos(60)}
Hence the answer is third opption\sqrt{3}F
Kiran Chauhan
4 years ago
As the charges are same. So they will repel each other. 
Here the triangle is equilateral. So the angle made will be 60 degree. 
Hence net force=√f2+f2+2×f×fcos60
=√2f2+2f2×1/2
=√3f2
=√3×f 
 
Hence root 3×f is the answer.